Write a 3rd degree polynomial function that has the given zeros.
-4,3,-1

Answer by ikleyn(52946)

.
Write a 3rd degree polynomial function that has the given zeros.
-4,3,-1
------------------------------------------------------------

(x-(-4)*(x-3)*(x-(-1) = (x+4)*(x-3)*(x+1) = . . .

Open parentheses yourself if you want or if you need.